# Hilbert Space Fragmentation and Subspace Scar Time-Crystallinity in Driven Homogeneous Central-Spin Models

### v1

@@ III.1 Scar time-crystallinity in Ising, XX, XXX, and XXZ central-spin models
Figure 4 shows the staggered magnetization as a function of time (in terms of Floquet periods) for four different types of interactions and for initial states $\mathinner{|{J\downarrow}\rangle}$ and $\mathinner{|{(J-4)\downarrow}\rangle}$, with $N=2J=21$ satellite spins. For Ising interactions, both initial states show a robust time-crystalline behavior that survives for at least a million Floquet periods [Fig. 4(a)]. On the other hand, for Heisenberg (XXX) and XXZ-type interactions, time-crystallinity is stable for $\mathinner{|{J\downarrow}\rangle}$ but not for $\mathinner{|{(J-4)\downarrow}\rangle}$ on long time scales [Fig. 4(c,d)]. For XX interactions and in the absence of a Zeeman field, time-crystalline behavior persists for a much longer time for $\mathinner{|{J\downarrow}\rangle}$ compared to $\mathinner{|{(J-4)\downarrow}\rangle}$ [Fig. 4(b)], although it does eventually decay after $10^{5}$ periods. (We set the Zeeman field to zero in this case because it provides more stable time-crystallinity, as we show below.) Thus, an eigenstate of $\mathrm{I}_{t}^{z}$ with extreme eigenvalues (highly polarized states) show more robust time-crystalline behavior compared to less polarized eigenstates regardless of the type of interaction. We can understand this phenomenon based on the stability of HSF discussed in section II.2.
@@
Recently, a conventional (MBL-type) time-crystalline behavior was shown to occur in Heisenberg and XXZ-type central-spin models with random coupling disorder for all initial product states [50]. In that work, the ergodicity breaking was likely related to MBL since random coupling disorder was necessary to achieve stable non-ergodicity. Additionally, the emergence of time-crystallinity required either a large Zeeman energy difference between the central and satellite spins or the application of additional pulses on the central spin every Floquet period. In contrast, here the time-crystalline behavior is due to stable HSF for fully polarized satellite states that have high overlap with Floquet scar states. Therefore, here time-crystallinity does not require random coupling disorder but only a large Zeeman energy difference between central and satellite spins, and it arises for only a few initial states. Interestingly, we see in Fig. 4(b) that in the case of the XX central-spin model, a Zeeman energy difference is not necessary for time-crystallinity for the initial state $\mathinner{|{J\downarrow}\rangle}$. We further examine the behavior of the XX model under random coupling disorder (Gaussian distribution) for $B_{z}=0$ in Appendix D, where we find that as we increase the disorder in the coupling terms, the stability of the time-crystallinity decreases. Thus, the XX central-spin model does not show the conventional (MBL type) time-crystalline behavior, but does exhibit scar time-crystallinity. We further discuss the nature of scar time-crystallinity in Sec. III.3. In the next subsection, we will see that for more general XXZ interactions, we can also observe scar time-crystallinity without a Zeeman energy difference.
III.2 Phase diagrams
@@
Figure 6: Non-equilibrium phase diagrams: Staggered satellite total spin time-averaged over $N_{C}=10^{4}$ Floquet cycles (Eq. (16)) as a function of the pulse error ($\theta$) and interaction strength ($A_{z}$ or $A_{xy}$ or $A=A_{z}=A_{xy}$). Phase diagrams are shown for (a) Ising, (b) XX, (c) Heisenberg, (d) XXZ interactions. The parameters are $B_{z}=100$ MHz, $\omega=1$ MHz. In (d) $A_{xy}=4.5$ MHz. The initial state is $\mathinner{|{J\uparrow}\rangle}$ and $N=2J=21$ in all cases.
@@
In Fig. 6, we show a different set of phase diagrams in which we consider the dependence of the staggered magnetization on the pulse error ($\theta$) instead of the Zeeman field. Overall, we see that Ising interactions are much more robust against the pulse error [Fig. 6 (a)] especially when $A_{z}/\omega$ is an odd integer. This is because, for these values of $A_{z}/\omega$, the first-order perturbative term vanishes in the evolved state and, hence, in $\mathinner{\langle{\psi(2pT,\theta)}|{\mathrm{I}_{t}^{z}}|{\psi(2pT,\theta)}\rangle}$ (see Appendix B.1 for details). While Heisenberg interactions show a similar robustness only for smaller values of the coupling strength [Fig. 6(c)]. In the case of XX interactions [Fig. 6 (b)], the system shows ergodic behavior for all considered values of $A_{xy}$ when $\theta>0.1\pi$; the narrow phase regions that extend to high values of $\theta$ in the case of Ising or Heisenberg interactions are no longer evident here. Anisotropic XXZ interactions [Fig. 6 (d)], on the other hand, produce phase regions that resemble those of the XX model for lower values of $A_{z}$, while they become more like those of the Ising model for larger values of $A_{z}$. This is as expected since the model interpolates between the Ising and XX models as $A_{z}$ is swept from zero up to values large compared to $A_{xy}$.
@@
Central-spin interactions
MBL-DTC (random couplings)
Scar-DTC (homogeneous couplings)
Remarks
No interaction
No
No
No DTC without interaction.
Ising
Yes
Yes
DTC is robust except when $A_{z}$ is an even integer multiple of $\omega$.
XX
No
Yes
Robust scar DTC exists for $B_{z}=0$, but it is not stable at large $B_{z}$.
Heisenberg (XXX)
Yes
Yes
Robust scar and MBL DTCs exist for large Zeeman energy mismatch between central and satellite spins.
XXZ
Yes
Yes
Robust scar DTC exists, but MBL DTC only exists for large Zeeman energy mismatch between central and satellite spins.
Table 1: Summary of when MBL and scar DTCs can occur in central-spin systems depending on the type of interaction and other factors. MBL DTCs require disorder in the couplings, while scar DTCs do not.
III.3 Scar versus conventional discrete time crystals
@@
Regarding point (ii), we saw in Sec. II.2 that for interactions of non-Ising type (XX, XXX, XXZ) and for most initial states of the form $\mathinner{|{m_{j}\sigma}\rangle}$, there is a 4D Floquet-Krylov subspace that breaks the discrete-time translation symmetry. The eigenstates of $U_{F}$ within these subspaces again have a Schrödinger cat-like form, making them physically inaccessible. However, only subspaces that contain eigenstates of $\mathrm{I}_{t}^{z}$ with extreme eigenvalues (highly polarized states) show robustness against pulse errors (see Fig. 4(b,c,d)). We can think of the conventional DTC as a special case of this “subspace" scar DTC in which the dimension of each subspace is one, and where all initial states are scar states.
@@ IV Experimental implementations
A key experimental requirement for realizing DTCs in electron-nuclear central-spin systems is the ability to control both the electronic as well as the nuclear spins. Control of the electronic spin is routinely done in experiments using either microwave or optical pulses. Direct control of nuclear spins has also been demonstrated using separate RF fields, although this approach requires additional experimental overhead [56, 46]. Alternatively, one can instead drive nuclear spin rotations indirectly by using carefully designed pulse sequences that are applied to the electronic spin [57, 54, 58, 59, 55, 60]. Nuclear spin control is necessary for both initializing the system in a highly polarized state and for performing the $\pi$ rotations that are needed every Floquet period. Refs. [57, 46, 60, 54] demonstrated the initialization and pulsing of a small set of nuclear spins in the vicinity of color centers or rare-earth ions via indirect control through the electronic spin, in some cases with the assistance of an additional RF field. In the case of quantum dots, where the nuclei constitute a dense spin ensemble, individual addressing of nuclear spins is not possible. However, initialization of the nuclei into highly polarized states is still possible using indirect electron-mediated control [61, 59]. Moreover, control of individual nuclear magnon modes in these systems has also been demonstrated [41, 61, 40].
[…]
Overall, there are many promising avenues for the potential experimental realization of scar time-crystalline physics in electron- or hole-nuclear central-spin systems in a variety of platforms, with many of the key capabilities already demonstrated.
@@ V Conclusion and Outlook
We have shown that Hilbert space fragmentation and scar time-crystallinity can arise in various types of periodically driven central-spin models with homogeneous interactions. The dynamical formation of two-dimensional or four-dimensional Floquet-Krylov subspaces leads to the spontaneous breaking of discrete time-translation symmetry and gives rise to a subharmonic response in the total satellite spin magnetization. We further showed that this response persists in the presence of pulse errors only for highly polarized initial states—a defining feature of scar time-crystallinity. We mapped out the scar DTC phase regions of Ising, XX, Heisenberg, and XXZ central-spin models by computing the dependence of the time-averaged satellite spin magnetization on the applied magnetic field, the interaction strength, and the pulse error. Unlike conventional MBL-type DTCs, this type of scar time-crystallinity does not require coupling disorder and can occur in the absence of a Zeeman energy mismatch between central and satellite spins.
Our work introduces a new type of “subspace scar time crystal" in which a time-evolving subspace, rather than an individual quantum state, breaks the discrete time-translation symmetry of the Hamiltonian. In a recent theoretical work [62], the breaking of discrete time-translation symmetry by Floquet-Krylov (cellular automata) subspaces has been suggested, but not explicitly demonstrated in an example. Thus, our work demonstrates, to the best of our knowledge, the first stable subspace scar time-crystalline behavior without any disorder.
@@ B.1 Ising type Hamiltonian
$\sum_{\begin{subarray}{c}r=2q,\\ q\in\mathbb{N}\end{subarray}}^{2p-2}U_{F}^{r}PU_{F}^{2p-r}\mathinner{|{m\uparrow}\rangle}$ $=\sum_{\begin{subarray}{c}r=2q,\\ q\in\mathbb{N}\end{subarray}}^{2p-2}U_{F}^{r}\big{[}e^{-i(2p-r)A_{z}mT/2}\times$ (47)
@@
$\sum_{\begin{subarray}{c}r=2q-1,\\ q\in\mathbb{N}\end{subarray}}^{2p-1}U_{F}^{r}PU_{F}^{2p-r}\mathinner{|{m\uparrow}\rangle}=\sum_{\begin{subarray}{c}r=2q-1,\\ q\in\mathbb{N}\end{subarray}}^{2p-1}U_{F}^{r}\big{[}e^{-i(\Phi_{m}^{\uparrow}+\frac{2p-r-1}{2}A_{z}mT)}$ (53)
@@
Thus, using Eqs.  (B.1,47,53), we have up to ${\cal O}(\theta)$:,
@@
$\sum_{\begin{subarray}{c}r=2q,\\ q\in\mathbb{N}\end{subarray}}^{2p-2}e^{irA_{z}mT}+\sum_{\begin{subarray}{c}r=2q-1,\\ q\in\mathbb{N}\end{subarray}}^{2p-1}e^{i(rA_{z}mT-B_{z}T)}\}\mathinner{|{m\downarrow}\rangle}+$ (59)
@@
where $p\in\mathbb{Z}$. Thus from here, we see that when $A_{z}$ is an odd multiple of the drive frequency, we get a more robust time-crystalline behavior. In the main text, we considered $\mathinner{|{\psi(0)}\rangle}=\mathinner{|{J\uparrow}\rangle}$ in Sec. III.2. We also chose $J=N/2=10.5$ and integer values of $B_{z}/\omega$. In that case and for $A_{z}/\omega$ an odd integer, the coefficient of $\mathinner{|{m\downarrow}\rangle}=\mathinner{|{J\downarrow}\rangle}$ (as in Eq. (B.1)) is also zero. So, after $2p$ cycles, we obtain

### v2

@@ III.2 Scar time-crystallinity in Ising, XX, XXX, and XXZ central-spin models
Figure 4 shows the staggered magnetization as a function of time (in terms of Floquet periods) for four different types of interactions and for initial states $\mathinner{|{J\downarrow}\rangle}$ and $\mathinner{|{(J-4)\downarrow}\rangle}$, with $N=2J=21$ satellite spins. For Ising interactions, both initial states show a robust time-crystalline behavior that survives for at least a million Floquet periods [Fig. 4(a)]. On the other hand, for Heisenberg (XXX) and XXZ-type interactions, time-crystallinity is stable for $\mathinner{|{J\downarrow}\rangle}$ but not for $\mathinner{|{(J-4)\downarrow}\rangle}$ on long time scales [Fig. 4(c,d)]. For XX interactions and in the absence of a Zeeman field, time-crystalline behavior persists for a much longer time for $\mathinner{|{J\downarrow}\rangle}$ compared to $\mathinner{|{(J-4)\downarrow}\rangle}$ [Fig. 4(b)], although it does eventually decay after $10^{5}$ periods. (We set the Zeeman field to zero in this case because it provides more stable time-crystallinity, as we show below.) Thus, an eigenstate of $\mathrm{I}_{t}^{z}$ with extreme eigenvalues (highly polarized states) show more robust time-crystalline behavior compared to less polarized eigenstates regardless of the type of interaction. We can understand this phenomenon based on the stability of HSF discussed in section II.2. Furthermore, the discussion in Sections II.2 and II.3 highlights the weak breaking of ergodicity, suggesting that the time crystalline behavior for a fully polarized nuclear state, observed in Fig. 4, will persist for much longer than what is shown.
@@
Recently, a conventional (MBL-type) time-crystalline behavior was shown to occur in Heisenberg and XXZ-type central-spin models with random coupling disorder for all initial product states [54]. In that work, the ergodicity breaking was likely related to MBL since random coupling disorder was necessary to achieve stable non-ergodicity. Additionally, the emergence of time-crystallinity required either a large Zeeman energy difference between the central and satellite spins or the application of additional pulses on the central spin every Floquet period. In contrast, here the time-crystalline behavior is due to stable HSF for fully polarized satellite states that have high overlap with Floquet scar states. Therefore, here time-crystallinity does not require random coupling disorder but only a large Zeeman energy difference between central and satellite spins, and it arises for only a few initial states. Interestingly, we see in Fig. 4(b) that in the case of the XX central-spin model, a Zeeman energy difference is not necessary for time-crystallinity for the initial state $\mathinner{|{J\downarrow}\rangle}$. We further examine the behavior of the XX model under random coupling disorder (Gaussian distribution) for $B_{z}=0$ in Appendix D, where we find that as we increase the disorder in the coupling terms, the stability of the time-crystallinity decreases. Thus, the XX central-spin model does not show the conventional (MBL type) time-crystalline behavior, but does exhibit scar time-crystallinity. We further discuss the nature of scar time-crystallinity in Sec. III.4. In the next subsection, we will see that for more general XXZ interactions, we can also observe scar time-crystallinity without a Zeeman energy difference.
III.3 Phase diagrams
@@
Figure 6: Non-equilibrium phase diagrams: Staggered satellite total spin time-averaged over $N_{C}=10^{4}$ Floquet cycles (Eq. (17)) as a function of the pulse error ($\theta$) and interaction strength ($A_{z}$ or $A_{xy}$ or $A=A_{z}=A_{xy}$). Phase diagrams are shown for (a) Ising, (b) XX, (c) Heisenberg, (d) XXZ interactions. The parameters are $B_{z}=100$ MHz (in (a,c,d)), $B_{z}=0$ MHz (in (b)), $\omega=1$ MHz. In (d) $A_{xy}=4.5$ MHz. The initial state is $\mathinner{|{J\uparrow}\rangle}$ and $N=2J=21$ in all cases.
@@
In Fig. 6, we show a different set of phase diagrams in which we consider the dependence of the staggered magnetization on the pulse error ($\theta$) instead of the Zeeman field. Overall, we see that Ising interactions are much more robust against the pulse error [Fig. 6 (a)] especially when $A_{z}/\omega$ is an odd integer. This is because, for these values of $A_{z}/\omega$, the first-order perturbative term vanishes in the evolved state and, hence, in $\mathinner{\langle{\psi(2pT,\theta)}|{\mathrm{I}_{t}^{z}}|{\psi(2pT,\theta)}\rangle}$ (see Appendix B.1 for details). While Heisenberg interactions show a similar robustness only for smaller values of the coupling strength [Fig. 6(c)]. In the case of XX interactions without Zeeman field [Fig. 6 (b)], the system shows ergodic behavior for all considered values of $A_{xy}$ when $\theta>0.1\pi$; the narrow phase regions that extend to high values of $\theta$ in the case of Ising or Heisenberg interactions are no longer evident here. Anisotropic XXZ interactions [Fig. 6 (d)], on the other hand, produce phase regions that resemble those of the XX model for lower values of $A_{z}$, while they become more like those of the Ising model for larger values of $A_{z}$. This is as expected since the model interpolates between the Ising and XX models as $A_{z}$ is swept from zero up to values large compared to $A_{xy}$.
@@
Table 1: Summary of when MBL and scar DTCs occur in central-spin systems depending on the type of interaction and other parameters. MBL DTCs require disorder in the couplings, while scar DTCs do not.
Interaction Type MBL-DTC (Random Couplings) Scar-DTC (Homogeneous Couplings) Remarks
None No No DTC does not emerge without interactions.
Ising Yes Yes DTC is robust except at certain discrete values of $A_{z}/\omega$. Robustness is maximized around odd values of $A_{z}/\omega$.
XX No Yes A robust scar DTC exists at $B_{z}=0$, but it becomes unstable for large $B_{z}$.
Heisenberg (XXX) Yes Yes Robust MBL and scar DTCs emerge when there is a large Zeeman energy mismatch between the central and satellite spins.
XXZ Yes Yes A robust scar DTC exists. An MBL DTC emerges only when there is a large Zeeman energy mismatch between the central and satellite spins.
III.4 Scar versus conventional discrete time crystals
@@
Regarding point (ii), we saw in Sec. II.2 that for interactions of non-Ising type (XX, XXX, XXZ) and for most initial states of the form $\mathinner{|{m_{j}\sigma}\rangle}$, there is a 4D Floquet-Krylov subspace that breaks the discrete-time translation symmetry. The eigenstates of $U_{F}$ within these subspaces again have a Schrödinger cat-like form, making them physically inaccessible. However, only subspaces that contain eigenstates of $\mathrm{I}_{t}^{z}$ with extreme eigenvalues (highly polarized states) show robustness against pulse errors (see Fig. 4(b,c,d)). We can think of the conventional DTC as a special case of this “subspace” scar DTC in which the dimension of each subspace is one, and where all initial states are scar states.
@@ IV Experimental implementations
A key experimental requirement for realizing DTCs in electron-nuclear central-spin systems is the ability to control both the electronic as well as the nuclear spins. Control of the electronic spin is routinely done in experiments using either microwave or optical pulses. Direct control of nuclear spins has also been demonstrated using separate RF fields, although this approach requires additional experimental overhead [60, 50]. Alternatively, one can instead drive nuclear spin rotations indirectly by using carefully designed pulse sequences that are applied to the electronic spin [61, 58, 62, 63, 59, 64]. Nuclear spin control is necessary for both initializing the system in a highly polarized state and for performing the $\pi$ rotations that are needed every Floquet period. Refs. [61, 50, 64, 58] demonstrated the initialization and pulsing of a small set of nuclear spins in the vicinity of color centers or rare-earth ions via indirect control through the electronic spin, in some cases with the assistance of an additional RF field. Specifically, in weakly-coupled systems, dynamical decoupling sequences of $\pi$ pulses can be used to coherently interact an electron spin with a single nucleus while decoupling the rest of the system. These sequences enable coherent electron-nuclear dynamics, facilitating full nuclear control across various platforms [61, 50, 65]. Even though many-body dynamics may seem complex, dynamical decoupling can isolate an effective two-body system by decoupling other spins, maintaining the interaction between the electron spin and a single nuclear spin. Ideally, complete manipulation of the nuclear spin could flip the entire magnetization of the spin bath. In the case of quantum dots, where the nuclei constitute a dense spin ensemble, individual addressing of nuclear spins is not possible. However, initialization of the nuclei into highly polarized states is still possible using indirect electron-mediated control [66, 63]. Moreover, control of individual nuclear magnon modes in these systems has also been demonstrated [41, 66, 40].
Another important issue in realizing scar time-crystallinity in physical central-spin systems is the distribution of electron-nuclear couplings in these systems. In the case of color centers in diamond, for example, spinful 13C nuclei are randomly distributed throughout the diamond lattice with a natural abundance of 1.1%. The distance and orientation of each nucleus relative to the color center determines the magnitudes of the hyperfine couplings, and so significant coupling disorder is naturally present in the system. An effective coupling uniformity could be achieved dynamically by adjusting the spacing of pulses applied to the electronic spin in such a way as to select out only a few nuclei with similar couplings to participate in the joint electron-nuclear dynamics. This approach essentially amounts to a dynamical decoupling of all other nuclei from the color center. In the case of rare-earth ion systems such as a 171Yb3+ ion in YVO4 that couples to nearby 51V5+ nuclei, the 99.75% natural abundance of the latter ensures that the nuclear spins naturally form sets that uniformly couple to the 171Yb3+ electronic spin [58]. The situation is somewhat more complex in quantum dots, where electron-nuclear hyperfine couplings form a dense, Gaussian-like distribution. In this case, one could imagine using dynamical decoupling techniques on the electronic spin to single out annular subsets of nuclei with equal couplings, while decoupling all other nuclear spins.
Overall, there are many promising avenues for the potential experimental realization of scar time-crystalline physics in electron- or hole-nuclear central-spin systems in a variety of platforms, with many of the key capabilities already demonstrated. We also recognize that combining all of these key capabilities may still be challenging, but we are optimistic that this can be achieved in the future.
@@ V Conclusion and Outlook
We have shown that Hilbert space fragmentation and (subspace) scar time-crystallinity can arise in various types of periodically driven central-spin models with homogeneous interactions. The dynamical formation of two-dimensional or four-dimensional Floquet-Krylov subspaces leads to the spontaneous breaking of discrete time-translation symmetry and gives rise to a subharmonic response in the total satellite spin magnetization. We further showed that this response persists in the presence of pulse errors only for highly polarized initial states—a defining feature of scar time-crystallinity. We mapped out the scar DTC phase regions of Ising, XX, Heisenberg, and XXZ central-spin models by computing the dependence of the time-averaged satellite spin magnetization on the applied magnetic field, the interaction strength, and the pulse error. Unlike conventional MBL-type DTCs, this type of scar time-crystallinity does not require coupling disorder and can occur in the absence of a Zeeman energy mismatch between central and satellite spins.
Our work also introduces a new type of TC called “subspace scar time crystal” in which a time-evolving subspace, rather than an individual quantum state, breaks the discrete time-translation symmetry of the Hamiltonian. In a recent theoretical work [67], the breaking of discrete time-translation symmetry by Floquet-Krylov (cellular automata) subspaces has been suggested, but not explicitly demonstrated in an example. Thus, our work demonstrates, to the best of our knowledge, the first stable subspace scar time-crystalline behavior without any disorder.
@@ B.1 Ising type Hamiltonian
$\sum_{\begin{subarray}{c}r=2q,\\ q\in\mathbb{N}\end{subarray}}^{2p-2}U_{F}^{r}PU_{F}^{2p-r}\mathinner{|{m\uparrow}\rangle}$ $=\sum_{\begin{subarray}{c}r=2q,\\ q\in\mathbb{N}\end{subarray}}^{2p-2}U_{F}^{r}\big{[}e^{-i(2p-r)A_{z}mT/2}\times$
@@
$\sum_{\begin{subarray}{c}r=2q-1,\\ q\in\mathbb{N}\end{subarray}}^{2p-1}U_{F}^{r}PU_{F}^{2p-r}\mathinner{|{m\uparrow}\rangle}=\sum_{\begin{subarray}{c}r=2q-1,\\ q\in\mathbb{N}\end{subarray}}^{2p-1}U_{F}^{r}\big{[}e^{-i(\Phi_{m}^{\uparrow}+\frac{2p-r-1}{2}A_{z}mT)}$
@@
Thus, using Eqs.  (B.1,B.1,B.1), we have up to ${\cal O}(\theta)$:,
@@
$\sum_{\begin{subarray}{c}r=2q,\\ q\in\mathbb{N}\end{subarray}}^{2p-2}e^{irA_{z}mT}+\sum_{\begin{subarray}{c}r=2q-1,\\ q\in\mathbb{N}\end{subarray}}^{2p-1}e^{i(rA_{z}mT-B_{z}T)}\}\mathinner{|{m\downarrow}\rangle}+$
@@
where $p\in\mathbb{Z}$. Thus from here, we see that when $A_{z}$ is an odd multiple of the drive frequency, we get a more robust time-crystalline behavior. In the main text, we considered $\mathinner{|{\psi(0)}\rangle}=\mathinner{|{J\uparrow}\rangle}$ in Sec. III.3. We also chose $J=N/2=10.5$ and integer values of $B_{z}/\omega$. In that case and for $A_{z}/\omega$ an odd integer, the coefficient of $\mathinner{|{m\downarrow}\rangle}=\mathinner{|{J\downarrow}\rangle}$ (as in Eq. (B.1)) is also zero. So, after $2p$ cycles, we obtain